# When ControlNet Meets Inexplicit Masks: A Case Study of ControlNet on its Contour-following Ability

### v2

@@ Abstract
ControlNet excels at creating content that closely matches precise contours in user-provided masks. However, when these masks contain noise, as a frequent occurrence with non-expert users, the output would include unwanted artifacts. This paper first highlights the crucial role of controlling the impact of these inexplicit masks with diverse deterioration levels through in-depth analysis. Subsequently, to enhance controllability with inexplicit masks, an advanced Shape-aware ControlNet consisting of a deterioration estimator and a shape-prior modulation block is devised. The deterioration estimator assesses the deterioration factor of the provided masks. Then this factor is utilized in the modulation block to adaptively modulate the model’s contour-following ability, which helps it dismiss the noise part in the inexplicit masks. Extensive experiments prove its effectiveness in encouraging ControlNet to interpret inaccurate spatial conditions robustly rather than blindly following the given contours. We showcase application scenarios like modifying shape priors and composable shape-controllable generation. Codes are available.
1 Introduction
@@
Figure 1: ControlNet tends to preserve contours for spatial controllable generation over multi-modal control inputs, where green denotes recalled contours and blue denotes missing ones. However, inexplicit masks cause catastrophic degradation of image fidelity and realism. This paper largely enhances its robustness in interpreting inexplicit masks with inaccurate contours.
@@
Text-to-Image (T2I) generation techniques Saharia et al. (2022); Ramesh et al. (2022, 2021) have greatly changed the content creation area with high-fidelity synthesized images. By generating content following user-provided guidance like contours and shapes, ControlNet Zhang et al. (2023a) stands out for its prominent capability of spatial control over T2I diffusion models and has become an indispensable tool for creation. The shape guidance always takes the format of segmentation masks, where the details of the contours and positions are perfectly reserved as shown in Fig. 1. However, such good contour-following ability [^2] The ability to preserve contours of the conditional inputs. of ControlNet may cause artifacts when there exists noise in the mask, especially for non-expert users that have difficulty in providing accurate masks. Unfortunately, its performance on inexplicit masks with inaccurate contours remains under-explored, and studies on how to employ inexplicit masks are ignored in current works. This issue hinders non-expert users from creating better images through ControlNet.
@@
To better assist non-expert users in generating satisfactory images, this paper focuses on utilizing inexplicit masks during the generation process. We first analyze the most important property, i.e., the contour-following ability, on the inexplicit mask-guided generation process through deteriorating masks and tuning hyperparameters. Specifically, we study this property quantitatively and reveal preliminary analyses on two key areas: 1) its performance on conditional masks at varying precision levels, and 2) the influence of hyperparameters on this property. One of the main findings is that masks with inaccurate contours would cause artifacts due to strong contour instructions. In other words, the contour-following ability implicitly assumes that the conditional inputs align with the shape priors of specific objects. Otherwise, it would cause artifacts or violate the spatial control in synthesized images as shown in Fig. 1. Given that precise conditional images are typically image-oriented (e.g., human annotation or those extracted from reference images by offline detectors) or expert-provided (e.g., artists), it constrains creation within the boundaries of existing images and experts. Obtaining precise control inputs can be cumbersome and challenging for non-experts. This fact largely restricts creation through scribbles, thus hindering broad applications of ControlNet.
@@
While an intuitive solution is to adjust hyperparameters to achieve satisfactory results, finding the optimal setting can be challenging. Besides, this strategy usually entails a trade-off between image fidelity and spatial control. Based on this observation, we improve ControlNet with the awareness of shape priors, namely Shape-aware ControlNet, which alleviates the impact of strong contour instructions and shows advances in robust interpretation of inexplicit masks. Specifically, our improvements comprise a deterioration estimator and a shape-prior modulation block to automatically adjust the model’s contour-following ability. The deterioration estimator assesses the deterioration ratio that depicts the similarity of the provided masks to explicit masks. Features from the Stable Diffusion (SD) Rombach et al. (2022) encoders are utilized during the estimation, as SD effectively encodes the shape priors of different objects. Then, we modulate this shape prior to the zero-convolution layers of ControlNet through the proposed shape-prior modulation block. This design enables extra control over the strength of contour instructions and adapts the model to the guidance of inaccurate contours. As Fig. 1 shows, our method follows object position and poses guided by inexplicit masks, keeping high fidelity and spatial control over T2I generation.
@@
This paper primarily focuses on ControlNet for two reasons. Firstly, ControlNet is the most representative and widely adopted method for spatially controllable T2I generation. It prevails over concurrent works for higher image quality and better preservation of outlines owing to the inherited powerful SD encoder and extensive training data. Secondly, ControlNet shares fundamental ideas and similar structures with existing adapter-based methods like T2I-Adapter Mou et al. (2023). Moreover, these methods exhibit similar properties (i.e., the contour-following ability) and face similar issues (i.e., performance degradation with inaccurate contours). Examples are provided in Appendix. Therefore, our conclusions and methods can be potentially extended to these methods.
@@
  • We study the contour-following ability of ControlNet quantitatively by examining its performance on masks of varying precision and hyperparameter settings. We reveal inexplicit masks would severely degrade image fidelity for strong priors induced by inaccurate contours.
@@
  • Our method adapts ControlNet to more flexible conditions like scribbles. We showcase its application scenarios like modifying object shapes and creative composable generation with masks of varying precision.
2 Related Works
@@
An intuitive solution to introduce spatial control is training models with spatially aligned conditions from scratch. Make-A-Scene Gafni et al. (2022) employs scene tokens derived from dense segmentation maps during generation, enabling complex scene generation and editing. SpaText Avrahami et al. (2023) extends it to open-vocabulary scenarios and introduces spatio-textual representations for sparse scene control. Composer Huang et al. (2023) decomposes images to representative factors like edges and then trains a model to recompose the input from these factors. However, training models from scratch requires large-scale training data and high computational costs.
@@
Another kind of work focuses on the utilization of adapters to inject spatial control into the pre-trained T2I models. For example, GLIGEN Li et al. (2023) involves gated self-attention layers to control the spatial layout during generation. T2I-adapter Mou et al. (2023) and ControlNet encode the spatial guidance via lightweight adapters or duplicated UNet structures, which are then fed into the decoder to control spatial structures like shapes and contours. Uni-ControlNet Zhao et al. (2023a), UniControl Qin et al. (2023), and Cocktail Hu et al. (2023) consolidate multi-modal conditions within a single framework. Since the base models are always frozen, these methods obtain better computational and data efficiency while retaining the generation ability of the base models. Despite these appealing characteristics, the impact of inexplicit masks is overlooked in these works, which may cause significant artifacts in generated images. Our proposed method can also be extended to these approaches to address this limitation.
3 Preliminary
@@
Figure 2: The metric curves of ControlNet-$\boldsymbol{m}_{r}$ on masks at varying levels of precision. The vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$, suffers from dramatic degradation on CLIP-Score and FID on deteriorated masks, but still keeps adhered to contours of provided masks.
ControlNet
@@
introduces a network to control T2I generation models, i.e., Stable Diffusion (SD), with extra spatially localized, task-specific conditional images including Canny edges, Hough lines, fake scribbles, key points, segmentation masks, shape normals, depths, and so on. Given one encoder block $F(\cdot;\theta)$ of SD parameterized by $\theta$, it accepts the input feature $\boldsymbol{x}$ and output $\boldsymbol{y}=F(\boldsymbol{x};\theta)$. ControlNet freezes the block and makes a trainable copy $F(\cdot;\theta^{\prime})$ with parameter $\theta^{\prime}$ to inject additional condition $\boldsymbol{c}$ with zero convolution, formulated as,
$\boldsymbol{y_{c}}=F(\boldsymbol{x};\theta)+\lambda*Z(F(\boldsymbol{x}+Z(\boldsymbol{c};w_{1});\theta^{\prime});w_{2}),$ (1)
@@
where $Z(\cdot;w_{1})$ and $Z(\cdot;w_{2})$ are two $1\times 1$ zero convolution layers with parameters $w_{1},w_{2}$ initialized with zeros for stable training. $\boldsymbol{y_{c}}$ is the output feature modulated by spatial control signals. $\lambda$, namely conditioning scale, is introduced during inference to adjust condition strength. Since features are additive, it allows compositions of multiple conditions, also known as the Multi-ControlNet Zhang et al. (2023a).
Classifier-free Guidance (CFG)
@@
Ho and Salimans (2022) is a widely employed technique in generative diffusion models to increase image quality. Given a latent noise $\boldsymbol{z}$, CFG is conducted through mixing samples generated by an unconditional model $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and a jointly trained conditional model $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$, which is computed as,
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z},\boldsymbol{c})=\epsilon_{\theta}(\boldsymbol{z},\boldsymbol{c})+\omega*(\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z},\boldsymbol{c})-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z})),$ (2)
@@
where $\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z},\boldsymbol{c})$ represents the final output under condition $\boldsymbol{c}$. $\omega$ is a scaling factor for a trade-off between image quality and diversity, namely CFG scale. In this paper, spatial conditions are injected into both $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$ when conducting classifier-free guidance with ControlNet.
4 Visiting Contour-following Ability
@@
This section studies the contour-following ability from two aspects, including the performance on masks of varying precision and its interaction with hyperparameters, i.e., CFG scale, conditioning scale, and condition injection strategy.
4.1 Setup
@@
We utilize the most common real-world control signal, i.e., object mask for experiments. The object mask is simple but effective in covering shape information. For simplicity, we focus on plain binary masks in experiments, which are the most straightforward signals in applications. This choice also helps to exclude the impacts of color and occlusion.
@@
Following Zhang et al. (2023b), our experiments involve COCO images Lin et al. (2014) and corresponding captions Chen et al. (2015). We utilize instance masks from LVIS Gupta et al. (2019), which offers sparse and precise human-annotated object masks of COCO images with over 1200 object classes. We filter out images containing empty annotations, resulting in 114k image-caption-mask triplets for training and 4.7k for testing, namely COCO-LVIS.
@@
We adopt SD v1.5 Runway (2022) as the base model, and ControlNet is trained from scratch on COCO-LVIS for ten epochs with a learning rate of $51e-$ 5. We take 50% prompt dropping for classifier-free guidance while keeping the SD parameters frozen. We use UniPC Zhao et al. (2023b) sampler with 50 sampling steps. To evaluate, we take human-annotated masks for control and generate four images per caption. CFG scale and conditioning scale are set to $7.5$ and $1.0$ if there is no extra illustration.
@@
CLIP-Score (ViT-L/14) Hessel et al. (2021) and FID Heusel et al. (2017) are adopted as two basic metrics to measure text-image alignment and image fidelity. In addition, as ControlNet tends to preserve all contour structures provided in the conditional images, we calculate the number of edge pixels retained in the generated images, namely Contour-Recall (CR), to measure the contour-following effects quantitatively. Supposing a conditional image $\boldsymbol{c}$ and generated images $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{N}$, CR is defined as the recall of edge pixels in the control inputs, formulated as,
$CR=\frac{1}{N}\sum_{i=1}^{N}\frac{|MaxPool(D(\boldsymbol{x}_{i}),\sigma)\cap(D(\boldsymbol{c}))|}{|D(\boldsymbol{c})|},$ (3)
@@
where $D(\cdot)$ is an edge detector that returns binary edge maps. $MaxPool(\cdot,\sigma)$ is a max-pooling function to tolerate an edge detection error of $\sigma$-pixels. In our implementation, we utilize the Canny edge detector and set a tolerance as $\sigma=2$. Higher CR indicates stronger contour-following ability.
4.2 Impact of Inexplicit Masks

### v3

@@ Abstract.
ControlNet excels at creating content that closely matches precise contours in user-provided masks. However, when these masks contain noise, as a frequent occurrence with non-expert users, the output would include unwanted artifacts. This paper first highlights the crucial role of controlling the impact of these inexplicit masks with diverse deterioration levels through in-depth analysis. Subsequently, to enhance controllability with inexplicit masks, an advanced Shape-aware ControlNet consisting of a deterioration estimator and a shape-prior modulation block is devised. The deterioration estimator assesses the deterioration factor of the provided masks. Then this factor is utilized in the modulation block to adaptively modulate the model’s contour-following ability, which helps it dismiss the noise part in the inexplicit masks. Extensive experiments prove its effectiveness in encouraging ControlNet to interpret inaccurate spatial conditions robustly rather than blindly following the given contours, suitable for diverse kinds of conditions. We showcase application scenarios like modifying shape priors and composable shape-controllable generation. Codes are available at github.
Text-to-Image Generation; ControlNet; Inexplicit Conditions; Shape-controllable Generation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680692ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Figure 1. ControlNet tends to preserve contours for spatial controllable generation over multi-modal control inputs, where green denotes recalled contours and blue denotes missing ones. However, inexplicit masks cause catastrophic degradation of image fidelity and realism. This paper largely enhances its robustness in interpreting inexplicit masks with inaccurate contours.
@@
Text-to-Image (T2I) generation techniques (Saharia et al., 2022; Ramesh et al., 2022, 2021; Esser et al., 2024; Pernias et al., 2023; Sauer et al., 2023) have greatly changed the content creation area with high-fidelity synthesized images. By generating content following user-provided guidance like contours and shapes, ControlNet (Zhang et al., 2023a) stands out for its prominent capability of spatial control over T2I diffusion models and has become an indispensable tool for creation. The shape guidance always takes the format of segmentation masks, where the details of the contours and positions are perfectly reserved as shown in Fig. 1. However, such good contour-following ability [^1] The contour-following ability refers to the ability to preserve the contours of conditional inputs in this paper. of ControlNet may cause artifacts when there exists noise in the mask, especially for non-expert users that have difficulty in providing accurate masks. Unfortunately, its performance on inexplicit masks with inaccurate contours remains under-explored, and studies on how to employ inexplicit masks are ignored in current works. This issue hinders non-expert users from creating better images through ControlNet.
@@
To better assist non-expert users in generating satisfactory images, this paper focuses on utilizing inexplicit masks during the generation process. We first analyze the most important property, i.e., the contour-following ability, on the inexplicit mask-guided generation process through deteriorating masks and tuning hyperparameters. Specifically, we study this property quantitatively and reveal preliminary analyses on two key areas: 1) its performance on conditional masks of varying deterioration degrees, and 2) the influence of hyperparameters on this property. One of the main findings is that masks with inaccurate contours would cause artifacts due to strong contour instructions (§ 4.2). In other words, the contour-following ability implicitly assumes that the conditional inputs align with the shape priors of specific objects. Otherwise, it would cause artifacts or violate the spatial control in synthesized images as shown in Fig. 1. Given that precise conditional images are typically image-oriented (e.g., human annotation or those extracted from reference images by offline detectors) or expert-provided (e.g., artists), it constrains creation within the boundaries of existing images and experts. However, obtaining precise control inputs can be cumbersome and challenging for non-experts. This fact largely restricts creation through human scribbles, thus hindering broad applications of ControlNet.
@@
While an intuitive solution is to adjust hyperparameters to achieve satisfactory results, finding the optimal setting can be challenging (§ 4.3). Besides, this strategy usually entails a trade-off between image fidelity and spatial control. Based on this observation, we improve ControlNet with the awareness of shape priors, namely Shape-aware ControlNet (§ 5), which alleviates the impact of strong contour instructions and shows advances in robust interpretation of inexplicit masks (§ 6). Specifically, our improvements comprise a deterioration estimator and a shape-prior modulation block to automatically adjust the model’s contour-following ability. The deterioration estimator assesses the deterioration ratio that depicts the similarity of the provided masks to explicit masks. Features from the Stable Diffusion (SD) (Rombach et al., 2022) encoders are utilized during the estimation, as SD effectively encodes the shape priors of different objects. Then, we modulate this shape prior to the zero-convolution layers of ControlNet through the proposed shape-prior modulation block. This design enables extra control over the strength of contour instructions and adapts the model to the guidance of inaccurate contours. As Fig. 1 shows, our method follows object position and poses guided by inexplicit masks, keeping high fidelity and spatial control over T2I generation. It also adapts to other conditions beyond masks (§ 6.4). Moreover, though we only employ dilated masks for training, our method generalizes well to more realistic masks including user scribbles and programmatic TikZ sketches (§ 6.5).
@@
This paper primarily focuses on ControlNet for two reasons. Firstly, ControlNet is the most representative and widely adopted method for spatially controllable T2I generation. It prevails over concurrent works for higher image quality and better preservation of outlines owing to the inherited powerful SD encoder and extensive training data. ControlNet is influential and has been widely incorporated in tasks including image animation (Wang et al., 2023; Xu et al., 2023), video generation (Guo et al., 2023; Liang et al., 2023), and 3D generation (Chen et al., 2023), etc. Secondly, ControlNet shares fundamental ideas and similar structures with existing adapter-based methods like T2I-Adapter (Mou et al., 2024). Moreover, these methods exhibit similar properties (i.e., the contour-following ability) and face similar issues (i.e., performance degradation with inaccurate contours). More examples of ControlNet and T2I-Adapter are provided in Appendix Fig. S6. Therefore, our conclusions and methods can be potentially extended to these methods.
@@
  • We study the contour-following ability of ControlNet quantitatively by examining its performance on deteriorated masks of varying degrees under different hyperparameter settings. We reveal inexplicit masks would severely degrade image fidelity for strong shape priors induced by inaccurate contours.
@@
  • Our method adapts ControlNet to more flexible conditions like scribbles, sketches, and more condition types beyond masks. We showcase its application scenarios like modifying object shapes and creative composable generation with deteriorated masks of varying degrees.
2. Related Works
@@
An intuitive solution to introduce spatial control is training models with spatially aligned conditions from scratch. Make-A-Scene (Gafni et al., 2022) employs scene tokens derived from dense segmentation maps during generation, enabling complex scene generation and editing. SpaText (Avrahami et al., 2023) extends it to open-vocabulary scenarios and introduces spatio-textual representations for sparse scene control. Composer (Huang et al., 2023) decomposes images to representative factors like edges and then trains a model to recompose the input from these factors. However, training models from scratch requires large-scale training data and high computational costs.
@@
Another kind of work focuses on the utilization of adapters to inject spatial control into the pre-trained T2I models. For example, GLIGEN (Li et al., 2023) involves gated self-attention layers to control the spatial layout during generation. T2I-adapter (Mou et al., 2024) and ControlNet encode the spatial guidance via lightweight adapters or duplicated UNet encoders, which are then fed into the decoder to inject spatial structures like shapes and contours. Uni-ControlNet (Zhao et al., 2024), UniControl (Qin et al., 2024), and Cocktail (Hu et al., 2023) consolidate multi-modal conditions within a single framework. Since the base models are always frozen, these methods obtain better computational and data efficiency while retaining the generation ability of the base models. Despite these appealing characteristics, the impact of inexplicit masks is overlooked in these works, which may cause significant artifacts in generated images. Our proposed method can also be extended to these approaches to address this limitation.
3. Preliminary
@@
Figure 2. The metric curves of ControlNet-$\boldsymbol{m}_{r}$ on masks of varying deterioration degrees. The vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$, suffers from dramatic degradation on CLIP-Score and FID on deteriorated masks, but keeps adhered to the contours of provided masks. ControlNet-$\boldsymbol{m}_{r}$ ($r>0$) exhibits more robust performance on deteriorated masks as the dilation radius $r$ becomes larger.
@@
ControlNet (Zhang et al., 2023a) introduces a network as the adapter to control T2I generation models, i.e., Stable Diffusion (SD), with extra spatially localized, task-specific conditional images including edges (Xie and Tu, 2015; Xuan et al., 2023), Hough lines (Gu et al., 2022), fake scribbles (Xie and Tu, 2015), key points (Cao et al., 2019), segmentation masks, shape normals, depths (Ranftl et al., 2022), and so on. Supposing one encoder block $F(\cdot;\theta)$ of SD parameterized by $\theta$, it accepts the input feature $\boldsymbol{x}$ and output $\boldsymbol{y}=F(\boldsymbol{x};\theta)$. ControlNet freezes the block and makes a trainable copy $F(\cdot;\theta^{\prime})$ with parameter $\theta^{\prime}$ to inject additional condition $\boldsymbol{c}$ with zero convolution, formulated as,
(1) $\boldsymbol{y_{c}}=F(\boldsymbol{x};\theta)+\lambda*Z(F(\boldsymbol{x}+Z(\boldsymbol{c};w_{1});\theta^{\prime});w_{2}),$
@@
where $Z(\cdot;w_{1})$ and $Z(\cdot;w_{2})$ are two $1\times 1$ zero convolution layers with parameters $w_{1},w_{2}$ initialized with zeros for stable training. $\boldsymbol{y_{c}}$ is the output feature modulated by spatial control signals. $\lambda$, namely conditioning scale, is introduced during inference to adjust condition strength. Since features are additive, it allows compositions of multiple conditions, also known as Multi-ControlNet (Zhang et al., 2023a).
@@
Classifier-free Guidance (CFG) (Ho and Salimans, 2021) is a widely employed technique in generative diffusion models to increase image quality. Given a latent noise $\boldsymbol{z}$, CFG is conducted through mixing samples generated by an unconditional model $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and a jointly trained conditional model $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$, which is computed as,
(2) $\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z},\boldsymbol{c})=\epsilon_{\theta}(\boldsymbol{z},\boldsymbol{c})+\omega*(\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z},\boldsymbol{c})-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z})),$
@@
where $\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z},\boldsymbol{c})$ denotes the final output under condition $\boldsymbol{c}$. $\omega$ is a scaling factor for a trade-off between image quality and diversity, namely CFG scale. In this paper, the spatial conditions are injected into both $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$ when conducting CFG with ControlNet.
4. Visiting the Contour-following Ability
@@
This section studies the contour-following ability from two aspects, including the performance on deteriorated masks of varying degrees and its interaction with hyperparameters, i.e., CFG scale, conditioning scale, and condition injection strategy.
4.1. Setup
@@
We utilize the most common real-world control signal, i.e., object mask for experiments. The object mask is simple but effective in covering shape information. For simplicity, we focus on plain binary masks in experiments, which are the most straightforward user-provided signals in real-world applications. This choice also helps to exclude the impacts of color and occlusion.
@@
Following Control-GPT (Zhang et al., 2023b), our experiments involve COCO images (Lin et al., 2014) and corresponding captions (Chen et al., 2015). We utilize instance masks from LVIS (Gupta et al., 2019), which offers sparse and precise human-annotated object masks of COCO images with over 1,200 object classes. We filter out images containing empty annotations, resulting in 114k image-caption-mask triplets for training and 4.7k for testing, namely the COCO-LVIS dataset.
@@
We adopt SD v1.5 (Runway, 2022) as the base model, and ControlNet is trained from scratch on COCO-LVIS for ten epochs with a learning rate of $51e-$ 5 for all experiments. We take 50% prompt dropping for classifier-free guidance while keeping the SD parameters frozen. We use UniPC (Zhao et al., 2023) sampler with 50 sampling steps. To evaluate, we take ground-truth masks for control and generate four images per caption. CFG scale and conditioning scale are set to $7.5$ and $1.0$ if there is no extra illustration.
@@
For evaluation, CLIP-Score (ViT-L/14) (Hessel et al., 2021) and FID (Heusel et al., 2017) are adopted as two basic metrics to measure the text-image alignment and image fidelity. In addition, as ControlNet tends to preserve all contour structures provided in the conditional images, we calculate the number of edge pixels retained in the generated images, namely Contour-Recall (CR), to measure the contour-following effects quantitatively. Supposing a conditional image $\boldsymbol{c}$ and generated images $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{N}$, CR is defined as the recall of edge pixels in the control inputs, formulated as,
(3) $CR=\frac{1}{N}\sum_{i=1}^{N}\frac{|MaxPool(D(\boldsymbol{x}_{i}),\sigma)\cap(D(\boldsymbol{c}))|}{|D(\boldsymbol{c})|},$
@@
where $D(\cdot)$ is an edge detector that returns binary edge maps. We employ a max-pooling function $MaxPool(\cdot,\sigma)$ to tolerate an edge detection error of $\sigma$-pixels. In our implementation, we utilize the Canny edge detector as $D(\cdot)$ and set a tolerance as $\sigma=2$. A higher CR score indicates stronger contour-following ability.
4.2. Impact of Inexplicit Masks